Raleigh Company manufactures two joint products. At the split-off point, they have sales values of:

  • Product 1 $25 per unit
  • Product 2 $15 per unit

After further processing, the company can sell them for $35 and $22, respectively. Product 1 costs $12 per unit to process further and Product 2 costs $5 to process further. Should further processing be done on either or both of these products? (show and label calculations)
Answer #1

Answer- Product 2 should be further processed and Product 1 should be sell at split-off-point.

Explanation-

Raleigh Company
Statement Of Incremental Profit/(loss)
Product Sale Price at Split-Off-Point Sale Price of processed product Incremental sale per unit Cost per unit to further process Incremental profit/(loss) per unit
$ $ $ $ $
(a) (b) (c=b-a) (d) (e=c-d)
1 $25 per unit $35 per unit $10 per unit $12 per unit -2
2 $15 pe unit $22 per unit $7 per unit $5 per unit 2